OpenStudy (anonymous):

A job applicant estimates that his chance of passing a qualifying examination is 3/4, and his chance of being appointed if he does pass is 3/8 . What is the probability that he will receive the job? 1/2 9/32 3/8

OpenStudy (anonymous):

@iGreen I chose 3/8 and and I got it by subtracting 3/4-3/8. Is it correct Green?

OpenStudy (igreen):

I think you have to multiply them.. \(\dfrac{3}{4} \times \dfrac{3}{8} \rightarrow \dfrac{3 \times 3}{4 \times 8}\) Can you simplify that?

OpenStudy (igreen):

@VengefulWave

OpenStudy (anonymous):

@iGreen No I don't think so. Sorry had to reset my net.

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

Wait, I think it's 9/32 @iGreen